OpenStudy (anonymous):

Factor this polynomial expression: 2(2x-1)^2-9(2x-1)+4

jimthompson5910 (jim_thompson5910):

let z = 2x-1 which means you'll go from 2(2x-1)^2-9(2x-1)+4 to 2z^2-9z+4

jimthompson5910 (jim_thompson5910):

are you able to factor 2z^2-9z+4 ?

OpenStudy (anonymous):

I did that, and I'm normally excellent at this, but I'm having a mental block and can't factor that even with the z's.

jimthompson5910 (jim_thompson5910):

use the AC method A = 2 B = -9 C = 4 A*C = 2*4 = 8 find two numbers that multiply to 8 and also add to B = -9

jimthompson5910 (jim_thompson5910):

hint: here are all the ways to multiply two whole numbers (positive or negative) to get 8 1*8 = 8 2*4 = 8 (-1)*(-8) = 8 (-2)*(-4) = 8

OpenStudy (anonymous):

Thank you so much. YOU GET A MEDAL, YOU GET A MEDAL, YOU GET A MEDAL! EVERYONE GETS A MEDAL! https://youtu.be/WmCQ-V7c7Bc?t=3m

OpenStudy (anonymous):

@jim_thompson5910 I tried it and I got (4x-3)(2x-9) when the answer is supposed to be (4x-3)(2x-5)... I did 2z^2-9z+4, then I did the AC method to get (2z-1)(z-8). Then I replaced the z with the actual equation and got [2(2x-1)-1][2x-1-8]. I then simplified to (4x-3)(2x-9)...

jimthompson5910 (jim_thompson5910):

(2z-1)(z-8) isn't the correct factorization for 2z^2-9z+4

OpenStudy (anonymous):

2x4 is 8, so what is the correct factorization?

jimthompson5910 (jim_thompson5910):

2z^2-9z+4 2z^2-1z-8z+4 ... note: -1 and -8 add to -9 and multiply to +8 (2z^2-1z)+(-8z+4) z(2z-1)-4(2z-1) (z-4)(2z-1)

OpenStudy (anonymous):

Oh it's -2 and -4. But how am I supposed to know it's that and not -1 and -8 by just eyeballing it?

jimthompson5910 (jim_thompson5910):

(2z-1)(z-8) expands out to... (2z-1)(z-8) 2z*(z-8) - 1*(z-8) 2z^2 - 16z - z + 8 2z^2 - 17z + 8 ------------------------------------------------------- while (z-4)(2z-1) expands out to ... (z-4)(2z-1) z*(2z-1) - 4*(2z-1) 2z^2 - z - 8z + 4 2z^2 - 9z + 4

OpenStudy (anonymous):

Sorry I lost connection. Okay, I understand. Is there a way of eyeballing it to know it's not -1 and -8 without the calculation?

jimthompson5910 (jim_thompson5910):

Here's another way to factor 2z^2-9z+4 use the quadratic formula to solve 2z^2-9z+4 = 0 to get these two solutions: z = 4 or z = 1/2 use these two solutions to factor z = 4 or z = 1/2 z = 4 or 2z = 1 z-4 = 0 or 2z-1 = 0 (z-4)*(2z-1) = 0 notice how I'm working the zero product property in reverse

OpenStudy (anonymous):

Alright, thanks, but like I said is there a way of eyeballing what I said?

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean when you say "Is there a way of eyeballing it to know it's not -1 and -8 without the calculation?" the two numbers are -1 and -8. I don't know what you mean when you have "not" in there.

OpenStudy (anonymous):

The factor is -2 and -4 not -1 and -8, so I was wondering if there was a way of telling that it's the first and not the latter without expanding them both like you did a few comments ago?

jimthompson5910 (jim_thompson5910):

Factoring 2z^2-9z+4 using the AC method A = 2 B = -9 C = 4 A*C = 2*4 = 8 find two numbers that multiply to 8 and also add to B = -9 those two numbers are -1 and -8 (in either order). You arrive at this pair of numbers through guess and check. A list like you see below may help 1*8 = 8 2*4 = 8 (-1)*(-8) = 8 (-2)*(-4) = 8 ------------------------------------------------------- So again, the pair of numbers is -1 and -8. Use them to break up -9z and then factor by grouping 2z^2-9z+4 2z^2-1z-8z+4 (2z^2-1z)+(-8z+4) z(2z-1)-4(2z-1) (z-4)(2z-1)

jimthompson5910 (jim_thompson5910):

I'm still not sure why you're saying "The factor is -2 and -4 not -1 and -8"

OpenStudy (anonymous):

Somehow I skipped a step and it seemed way easier. Now it's more difficult but is correct! Thank you!